\begin{thebibliography}{99}

\bibitem{Alvareza} G. Alvareza, C. Senb, N. Furukawac, Y. Motomed and E. Dagottoe, The truncated polynomial expansion Monte Carlo method for fermion systems coupled to classical fieelds: a model independent implementation, Computer Physics Communications, 168(1)(2005), pp:32-45.

\bibitem{Awojoyogbe} O. B. Awojoyogbe and K. Boubaker. A solution to bloch NMR flow equations for the analysis of homodynamic functions of blood flow system using m-Boubaker polynomials, International  Journal of Current Applied Physics, Elsevier, DOI:10.1016/j.cap.2008.01.0193.

\bibitem{Bender} C.M. Bender and G. V. Dunne, Polynomials and operator orderings, Journal of Mathematical Physics 29(1988) pp: 1727-1731.

\bibitem{Boubaker1} K. Boubaker, On modifed Boubaker polynomials: some differential and analytical properties of the new polynomial issued from an attempt for solving bi-varied heat equation,Trends in Applied Sciences Research, 2(6)(2007) pp:540-544.

\bibitem{Chaouachi} A. Chaouachi, K. Boubaker, M. Amlouk and H. Bouzouita, Enhancement of pyrolysis spray disposal performance using thermal time-response to precursor uniform deposition. Eur.Phys. J. Appl. Phys. 37(2007)pp:105-109.

\bibitem{Choi} S-K. Choi, R. V. Grandhi, R. A. Canfeld and C. L. Pettit, Polynomial Chaos Expansion with Latin Hypercube Sampling for Estimating Response Variability , American Institute of Aeronautics and Astronautics (AIAA) Journal, 42 (2004),pp:1191-1198.

\bibitem{Guertz} B.J. Guertz, R. van Buuren and H. Lu, Application of polynomial preconditioners to conservation laws Application of polynomial preconditioners to conservation laws, Journal ofEngineering Mathematics 38 (2000) pp: 403-426.

\bibitem{Koelink} H. T. Koelink, The Addition Formula for Continuous q-Legendre Polynomials and Associated Spherical Elements on the SU(2) Quantum Group Related to AskeyCWilson Polynomials, SIAM Journal on Mathematical Analysis, 25(1)(1994), pp: 197-217.

\bibitem{Labiadh1}  H. Labiadh and K. Boubaker, A Sturm-Liouville shaped characteristic differential equation as a guide to establish a quasi-polynomial expression to the Boubaker polynomials, Differential Equationa and Control Processes, 2(2007), Electronic Journal,reg. No P2375 at 07.03.97 pp:117-133.

\bibitem{Labiadh2} H. Labiadh M. Dada, O.B. Awojoyogbe K. B. Ben Mahmoud and A. Bannour, Establishment of an ordinary generating function and a Christofel-Darboux type first-order differential equation for the heat equation related Boubaker-Turki polynomials, Journal of Differential Equations and C.P. 1(2008), 51-66.

\bibitem{Okada} R. Okada, N.Nakata, B. F. Spencer, K. Kasai and B.K. Saang, Rational polynomial approximation modeling for analysis of structures with VE dampers, Journal of Earthquake Engineering,10(2006), 97-125.

\bibitem{Philippou} A.N. Philippou, C. Georghiou, Convolutions of Fibonacci-type polynomials of order $k$ and the negative binomial distributions of the same order, Fibonacci Quart., 27 (1989), 209C216.

\bibitem{Luzon1} A. Luzon. Iterative processes related to Riordan arrays: The reciprocation and the inversion of power series. Preprint.

\bibitem{Luzon2} A. Luzon and M. A. Moron. Ultrametrics, Banach's fixed point theorem and the Riordan group. Discrete Appl. Math. 156 (2008) 2620-2635.

\bibitem{Luzon3} A. Luzon and M. A. Moron. Riordan matrices in the reciprocation of quadratic polynomials. Linear Algebra Appl. 430 (2009) 2254-2270.

\bibitem{Luzon4} A. Luzon and M. A. Moron. Recurrence relations for polynomial sequences via Riordan matrices. Preprint. See \url{http://arxiv.org/PS_cache/arxiv/pdf/0904/0904.2672v1.pdf}

\bibitem{Victor} Victor V. Prasolov Polynomials. Algorithms and Computation in Mathematics. Springer (2004).

\bibitem{Slama1} S. Slama, J. Bessrour, K. Boubaker and M.Bouhafs, Investigation of A3 point maximal front spatial evolution during resistance spot welding using 4q-Boubaker polynomial sequence, Proceedings of COTUME 2008, pp:79-80.

\bibitem{Sloan} Ian H. Sloan and R. S. Womersley, Good approximation on the sphere, with application to geodesy and the scattering of sound, Journal of Computational and App. Mathematics, 149 (2002), pp. 227-237

\bibitem{Zhao1} T. G. Zhao,B. K. Ben Mahmoud, M.A. Toumi, O. P. Faromika, M. Dada, O. B. Awojoyogbe, J. Magnuson, F. Lin, Some new properties of the applied-physics related Boubaker polynomials, Differential Equations and Control Processes,No1 (2009):7-19.

\bibitem{Zhao2} T. G. Zhao,Y. X. Wang,B. K. Ben Mahmoud, Limit and uniqueness of the Boubaker-Zhao polynomials single imaginary root sequence, International Journal of Mathematics and Computation, 1(8)(2008):13-16.

\bibitem{Bagula} Roger L. Bagula and Gary Adamson, Triangle of coefficients of Recursive Polynomials for Boubaker polynomials, OEIS (Encyclopedia of Integer Sequences), A137276 (2008).

\bibitem{Sloane} Neil J. A. Sloane, Triangle read by rows of  coefficients of Boubaker polynomial $B_n(x)$ in order of decreasing exponents, OEIS (Encyclopedia of Integer Sequences), A138034 (2008).

\bibitem{Ghanouchi1} J. Ghanouchi, H. Labiadh and  K. Boubaker, An attempt to solve the heat transfert equation in a model of pyrolysis spray using 4q-order m-Boubaker polynomials  International Journal of Heat and Technology, 26 (2008) pp. 49-53

\bibitem{Ghanouchi2} J. Ghanouchi, H. Labiadh and  K. Boubaker, An attempt to solve the heat transfert equation in a model of pyrolysis spray using 4q-order m-Boubaker polynomials  International Journal of Heat and Technology, 26 (2008) pp. 49-53

\bibitem{Slama2} S. Slama, J. Bessrour, K. Boubaker and M. Bouhafs, A dynamical model for investigation of A3 point maximal spatial evolution during resistance spot welding using Boubaker polynomials,  Eur. Phys. J. Appl. Phys. 44, (2008) pp. 317-322

\bibitem{Slama3} S. Slama, M. Bouhafs and  K. B. Ben Mahmoud,A Boubaker Polynomials Solution to Heat Equation for Monitoring A3 Point Evolution During Resistance Spot Welding,  International Journal of Heat and Technology,  26(2)  (2008) pp. 141-146.

\bibitem{Ghrib} T. Ghrib, K. Boubaker and  M. Bouhafs, Investigation of thermal diffusivity-microhardness  correlation  extended  to  surface-nitrured  steel using  Boubaker  polynomials  expansion, Modern Physics Letters B, 22, (2008) pp. 2893 - 2907

\bibitem{Boubaker2} K. Boubaker, The Boubaker polynomials, a new function class for solving bi-varied second order differential equations, F. E. Journal of App. Math. 31(2008) pp. 299 每 320.

\bibitem{Mahmoud1} B. K. Ben Mahmoud, Temperature 3D profiling in cryogenic cylindrical devices using Boubaker polynomials expansion scheme (BPES), Cryogenics,Volume 49, Issue 5(2009) pp. 217-220

\bibitem{Lazzez} S. Lazzez, K.B. Ben Mahmoud, S. Abroug, F. Saadallah, M. Amlouk, A Boubaker polynomials expansion scheme (BPES)-related protocol for measuring sprayed thin films thermal characteristics, Current Applied Physics Volume 9, Issue 5 (2009) pp.1129-1133

\bibitem{Fridjine1} S. Fridjine, K.B. Ben Mahmoud, M. Amlouk, M. Bouhafs, A study of sulfur/selenium substitution effects on physical and mechanical properties of vacuum-grown ZnS1?xSex compounds using Boubaker polynomials expansion scheme (BPES),Journal of Alloys and Compounds, Volume 479, Issues 1-2, (2009) pp. 457-461

\bibitem{Khelia} C. Kh\'{e}lia, K. Boubaker, T. Ben Nasrallah, M. Amlouk, S. Belgacem, Morphological and thermal properties of 汕-SnS2 sprayed thin films using Boubaker polynomials expansion, Journal of Alloys and Compounds, Volume 477, Issues 1-2, (2009) pp. 461-467

\bibitem{Mahmoud2} K.B. Ben Mahmoud, M. Amlouk, The $3D$ Amlouk每Boubaker expansivity每energy gap每Vickers hardness abacus: A new tool for optimizing semiconductor thin film materials,  Materials Letters, Volume 63, Issue 12 (2009) pp. 991-994

\bibitem{Dada}  M. Dada, O.B. Awojoyogbe, K. Boubaker, Heat transfer spray model: An improved theoretical thermal time-response to uniform layers deposit using Bessel and Boubaker polynomials, Current Applied Physics, Volume 9, Issue 3 (2009) pp. 622-624.

\bibitem{Rahmanov1} H. Rahmanov, Triangle read by rows: row $n$ gives coefficients of Boubaker polynomial $B_n(x)$, calculated for X=2cos(t), centered by adding -2cos(nt), then divided by 4, in order of decreasing exponents. OEIS (Encyclopedia of Integer Sequences), A160242

\bibitem{Rahmanov2} H. Rahmanov, Triangle read by rows: row $n$ gives values of the $4q-28$ Boubaker polynomials $B_{4q}(X)$ (named after Boubaker Boubaker (1897-1966)), calculated for $X=1$ (or -1). OEIS (Encyclopedia of Integer Sequences), A162180

\bibitem{Tabatabaei} S. Tabatabaei, T. Zhao, O. Awojoyogbe,  F. Moses, Cut-off cooling velocity profiling inside a keyhole model using the Boubaker polynomials expansion scheme,  Int.J. Heat Mass Transfer 45 (2009) pp. 1247-1255.

\bibitem{Fridjine2} S. Fridjine, M. Amlouk, A new parameter: An ABACUS for optimizig functional materials using the Boubaker polynomials expansion scheme, Modern Phys. Lett. B  23 (2009) pp. 2179-2182

\bibitem{Belhadj1} A. Belhadj, J. Bessrour, M. Bouhafs,  L. Barrallier,  Experimental and theoretical cooling velocity profile inside laser welded metals using keyhole approximation and Boubaker polynomials expansion, J. of Thermal Analysis and Calorimetry  97, (2009) pp. 911-920.

\bibitem{Belhadj2} A. Belhadj,  O. Onyango, N. Rozibaeva, Boubaker Polynomials Expansion Scheme-Related Heat Transfer Investigation Inside Keyhole Model , J. Thermophys. Heat Transf. 23 (2009) pp. 639-642.

\bibitem{Zhang} D. H. Zhang, F.W. Li, A Boubaker Polynomials Expansion Scheme BPES-Related Analytical  Solution to Williams-Brinkmann Stagnation Point Flow Equation at a Blunt Body,  Ir. Journal of App. Phys. Lett. IJAPLett.  2 (2009) pp. 25-30.

\end{thebibliography}

\documentclass[12pt]{article}

\usepackage[T1]{fontenc}
\usepackage[sc]{mathpazo}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{enumerate}
\usepackage{amsthm}
\usepackage{amsfonts,mathrsfs}

\usepackage{hyperref}
\hypersetup{pdfpagemode=UseNone}

\newenvironment{mylist}[1]{\begin{list}{}{
    \setlength{\leftmargin}{#1}
    \setlength{\rightmargin}{0mm}
    \setlength{\labelsep}{2mm}
    \setlength{\labelwidth}{8mm}
    \setlength{\itemsep}{0mm}}}
    {\end{list}}

\newcommand{\defeq}{\stackrel{\smash{\textnormal{\tiny def}}}{=}}

\newcommand{\Pa}[1]{\left(#1\right)}

\newcommand{\Br}[1]{\left[#1\right]}

\newtheorem{thrm}{Theorem}[section]

\theoremstyle{definition}
\newtheorem{definition}[thrm]{Definition}
\newtheorem{remark}[thrm]{Remark}

\numberwithin{equation}{section}

\newcounter{questionnumber}

\begin{document}

\vspace*{-8mm}


\begin{center}

{\Large\textbf{Fermat-linked relations for the Boubaker polynomial  sequences via Riordan matrices analysis}}\\[6mm]
{\large Karem Boubaker\footnote{E-mail: mmbb11112000@yahoo.fr}}\\
{\it\footnotesize ESSTT/ 63 Rue Sidi Jabeur 5100, Mahdia, Tunisia}\\[4mm]
{\large Lin Zhang\footnote{E-mail: godyalin@163.com}}\\
{\it\footnotesize Institute of Mathematics, Hangzhou Dianzi University,
Hangzhou, 310037, P.~R.~China}

\end{center}


\begin{abstract}

The Boubaker polynomials are investigated in this paper. Using Riordan matrices analysis, a sequence of relations outlining the relations with Chebyshev and Fermat polynomials have been obtained. The obtained expressions are a meaningful supply to recent applied physics studies using the Boubaker polynomials expansion scheme (BPES).\\~\\
\textbf{Keywords:} Riordan matrices; Boubaker polynomials\\~\\
\textbf{Mathematics Subject Classification 2000:} 33E20, 33E30, 35K05, 41A30, 41A55, 35K05, 33E30, 41A30, 41A55, 41A99.\\~\\
\textbf{PACS. 2008:} 02.30.Sa: Functional analysis - 47.10.A-: Mathematical formulations

\end{abstract}

\section{Introduction}

Polynomial expansion methods are extensively used in many mathematical and engineer fields to yield meaningful results for both numerical and analytical analysis \cite{Alvareza,Bender,Choi,Guertz,Koelink,Okada,Philippou,Sloan}. Among the most frequently used polynomials, the Boubaker polynomials are one of the interesting tools which were associated to several applied physics problems as well as the related polynomials such as the Boubaker-Turki polynomials \cite{Boubaker1,Bagula,Sloane,Ghanouchi1,Ghanouchi2,Slama2,Slama3,Ghrib,Boubaker2}, the $4-q$ Boubaker polynomials \cite{Zhao1} and the Boubaker-Zhao polynomials \cite{Zhao2}. For example, for some resolution purposes, a function $f(r)$ is expressed as an infinite nonlinear expansion of Boubaker-Zhao polynomials
\begin{eqnarray}
f(r) = \lim_{N\to+\infty}\Br{\frac{1}{2N}\sum^N_{n=1}\zeta_n \hat{B}_{4n}\Pa{r\frac{\alpha_n}{R}}},
\end{eqnarray}
where $\alpha_n$ are the minimal positive roots of the Boubaker $4n$-order polynomials $\hat{B}_{4n}$, $R$ is a maximum radial range and $\zeta_n$ are coefficients to be determined using the expression of $f(r)$. Since the Boubaker $4n$-order polynomials have the particular properties: for any $n$,
\begin{eqnarray}
\left\{\begin{array}{rcl}
         \hat{B}_{4n}(r)\left|\right._{r=0} &=& -2   \\
         \frac{\partial \hat{B}_{4n}(r)}{\partial r} &=& 0  \\
         \frac{\partial^2 \hat{B}_{4n}(r)}{\partial r^2} &=& 4n(n-1)
       \end{array}
\right.
\end{eqnarray}
The related the system (\ref{eq:1.3}) is induced:
\begin{eqnarray}\label{eq:1.3}
\left\{\begin{array}{rcl}
         f(0)&=& \lim_{N\to+\infty}\Br{\frac1{2N}\sum^N_{n=1}\zeta_n\hat{B}_{4n}\Pa{r\frac{\alpha_n}{R}}}\left|\right._{r=0} = -\frac1N\sum^N_{n=1}\zeta_n\\
         f(R)&=& \lim_{N\to+\infty}\Br{\frac1{2N}\sum^N_{n=1}\zeta_n\hat{B}_{4n}\Pa{r\frac{\alpha_n}{R}}}\left|\right._{r=R} = 0\\
         \frac{\partial f(r)}{\partial r}\left|\right._{r=0}&=& \lim_{N\to+\infty}\Br{\frac1{2N}\sum^N_{n=1}\zeta_n\frac{\partial\Pa{ \hat{B}_{4n}\Pa{r\frac{\alpha_n}{R}}}}{\partial r}}\left|\right._{r=0} = 0
       \end{array}
\right.
\end{eqnarray}

\section{The Boubaker polynomials}

The first monomial definition of the Boubaker polynomials \cite{Awojoyogbe,Boubaker1,Chaouachi,Labiadh1} appeared in a physical study that yielded an analytical solution to heat equation inside a physical model \cite{Labiadh2,Slama1}. This monomial definitions is traduced by (\ref{eq:2.1}):
\begin{definition}
A monomial definition of the Boubaker polynomials is:
\begin{eqnarray}\label{eq:2.1}
B_n(X) \defeq \sum^{\xi(n)}_{p=0} \Br{\frac{n-4p}{n-p}\binom{p}{n-p}}(-1)^p X^{n-2p},
\end{eqnarray}
where $\xi(n) = \lfloor\frac n 2\rfloor \defeq \frac{2n+(-1)^n -1}4$ (The symbol $\lfloor*\rfloor$ designates the floor function). Their coefficients could be defined through a recursive formula (\ref{eq:2.2}):
\begin{eqnarray}\label{eq:2.2}
\left\{\begin{array}{rcl}
         B_n(X) & = & \sum^{\xi(n)}_{j=0}\Br{b_{n,j}X^{n-2j}}, \\
         b_{n,0} & = & 1, \\
         b_{n,1} & = & -(n-4), \\
         b_{n,j+1} & = & \frac{(n-2j)(n-2j-1)}{(j+1)(n-j-1)}\cdot\frac{n-4j-4}{n-4j}\cdot b_{n,j}, \\
         b_{n,\xi(n)} & = & \left\{\begin{array}{rl}
                                     (-1)^{\frac n2}\cdot 2 & \text{if\ } n \text{\ even} \\
                                     (-1)^{\frac {n+1}2}\cdot (n-2) & \text{if\ } n \text{\ odd}
                                   \end{array}
         \right.
       \end{array}
\right.
\end{eqnarray}
\end{definition}

\begin{definition}
A recursive relation which yields the Boubaker polynomials is:
\begin{eqnarray}\label{eq:2.3}
\left\{\begin{array}{rcl}
         B_m(X) & = & XB_{m-1}(X) - B_{m-2}(X), \text{for\ } m>2,\\
         B_2(X) & = & X^2 + 2,\\
         B_1(X) & = & X,\\
         B_0(X) & = & 1.
       \end{array}
\right.
\end{eqnarray}
\end{definition}

\section{Riordan matrices of the Boubaker polynomials}

In this section, we will present Riordan matrices analysis of the Boubaker polynomials. The notations and the results of \cite{Luzon1,Luzon2,Luzon3,Luzon4} will be used extensively.
We start with the following relation (demonstrated on page 25 in \cite{Luzon4}):
\begin{eqnarray}\label{eq:3.1}
B_n(x) = U_n\Pa{\frac x2} + 3U_{n-2}\Pa{\frac x2},\quad \text{for\ } n\geqslant2
\end{eqnarray}
then:
\begin{eqnarray}
B_{2m}(x) &=& U_{2m}\Pa{\frac x2} + 3U_{2m-2}\Pa{\frac x2}\nonumber \\
&=& 2\sum^m_{k=0}\widetilde{T}_{2k}\Pa{\frac x2} + 6\sum^{m-1}_{k=0}\widetilde{T}_{2k}\Pa{\frac x2}\label{eq:3.2}\\
&=& 8\sum^{m-1}_{k=0}\widetilde{T}_{2k}\Pa{\frac x2} + 2\widetilde{T}_{2m}\Pa{\frac x2} = 4+ 8\sum^{m-1}_{k=0}T_{2k}\Pa{\frac x2} + 2T_{2m}\Pa{\frac x2}\label{eq:3.3}.
\end{eqnarray}
In a similar way:
\begin{eqnarray}
B_{2m+1}(x) &=& 8\sum^{m-1}_{k=0}\widetilde{T}_{2k+1}\Pa{\frac x2} + 2\widetilde{T}_{2m+1}\Pa{\frac x2} = 8\sum^{m-1}_{k=0}T_{2k+1}\Pa{\frac x2} + 2T_{2m+1}\Pa{\frac x2}\label{eq:3.4}\\
 &=& 8\sum^{m-1}_{k=0}\widetilde{T}_{2k}\Pa{\frac x2} + 2\widetilde{T}_{2m+1}\Pa{\frac x2}\label{eq:3.5}
\end{eqnarray}
so:
\begin{eqnarray}
B_{2m}(2\cos t) &=& 4+ 8\sum^{m-1}_{k=1}T_{2k}(\cos t) + 2T_{2m}(\cos t) \nonumber\\
&=& 4+ 8\sum^{m-1}_{k=1}\cos(2kt) + 2\cos(2mt)\label{eq:3.6}\\
B_{2m+1}(2\cos t) &=& 8\sum^{m-1}_{k=1}\cos((2k+1)t) + 2\cos((2m+1)t)\label{eq:3.7}.
\end{eqnarray}
Now, consider another new polynomial class defined by:
\begin{eqnarray}
B_n(2\cos t) &=& \frac{B_n(2\cos t) - 2T_n(\cos t)}4,\quad n>1\label{eq:3.8}
\end{eqnarray}
or:
\begin{eqnarray}\label{eq:3.9}
\left\{\begin{array}{ccc}
         B_n(x) & = & \frac{B_n(x)-2T_n\Pa{\frac x2}}4 \\
         x & = & 2\cos t
       \end{array}
\right.
\end{eqnarray}
So using Eq.~(\ref{eq:3.8}) and Eq.~(\ref{eq:3.9}) we get:
\begin{eqnarray}
B_{2m}(x) &=& \frac{B_{2m}(x) - 2T_{2m}\Pa{\frac x2}}4 = 1+ 2\sum^{m-1}_{k=0}T_{2k}\Pa{\frac x2},\label{eq:3.10}\\
B_{2m+1}(x) &=&\frac{B_{2m+1}(x) - 2T_{2m+1}\Pa{\frac x2}}4 = 2\sum^{m-1}_{k=0}T_{2k}\Pa{\frac x2}.\label{eq:3.11}
\end{eqnarray}
In order to obtain a generating function and to make a polynomial sequence (i. e. the degree is the subindex) we consider
$$
\widetilde{B}_n(x) = B_{n-2}(x).
$$
So, symbolically:
\begin{eqnarray}\label{eq:3.12}
\left[
  \begin{array}{c}
    \widetilde{B}_0(x) \\
    \widetilde{B}_1(x) \\
    \widetilde{B}_2(x) \\
    \widetilde{B}_3(x) \\
    \widetilde{B}_4(x) \\
    \widetilde{B}_5(x) \\
    M \\
  \end{array}
\right] = \left[
            \begin{array}{ccccccc}
              2& 0 & 0 & 0 & 0 & 0 & 0 \\
              0 & 2 & 0 & 0 & 0 & 0 & 0 \\
              2 & 0 & 2 & 0 & 0 & 0 & 0 \\
              0 & 2 & 0 & 2 & 0 & 0 & 0 \\
              2 & 0 & 2 & 0 & 2 & 0 & 0 \\
              0 & 2 & 0 & 2 & 0 & 2 & 0 \\
              M & M & M & M & M & M & O \\
            \end{array}
          \right]
\left[
  \begin{array}{c}
    \widetilde{T}_0(x) \\
    \widetilde{T}_1(x) \\
    \widetilde{T}_2(x) \\
    \widetilde{T}_3(x) \\
    \widetilde{T}_4(x) \\
    \widetilde{T}_5(x) \\
    M \\
  \end{array}
\right]
\end{eqnarray}
We can write this in terms of Riordan matrices in the next way:

\begin{eqnarray}\label{eq:3.13}
\sum_{n\geqslant0}\widetilde{B}_n(t) = T\Pa{\frac{2}{1-x^2}\big|1}T\Pa{\frac{1-x^2}4\big|\frac{1+x^2}2}T(2|2)\Pa{\frac1{1-tx}}.
\end{eqnarray}
or:
\begin{eqnarray}\label{eq:3.14}
\sum_{n\geqslant0}\widetilde{B}_n(t)x^n = T(1|1+x^2)\Pa{\frac1{1-tx}}.
\end{eqnarray}
In fact we have the Riordan matrix:
\begin{eqnarray}\label{eq:3.15}
T(1|1+x^2)
\end{eqnarray}
which is:
\begin{eqnarray}\label{eq:3.16}
\left[
  \begin{array}{ccccccc}
    1 & 0 & 0 & 0 & 0 & 0 & 0 \\
    0 & 1 & 0 & 0 & 0 & 0 & 0 \\
    -1 & 0 & 1 & 0 & 0 & 0 & 0 \\
    0 & -2 & 0 & 1 & 0 & 0 & 0 \\
    1 & 0 & -3 & 0 & 1 & 0& 0 \\
    0 & 3 & 0 & -4 & 0 & 1 & 0 \\
    M & M & M & M & M & M & O \\
  \end{array}
\right]
\end{eqnarray}
Hence, the few first $\widetilde{B}_n(x)$ are:
\begin{eqnarray}\label{eq:3.17}
\left\{\begin{array}{rcl}
         \widetilde{B}_0(x) & = & 1 \\
         \widetilde{B}_1(x) & = & x \\
         \widetilde{B}_2(x) & = & x^2-1 \\
         \widetilde{B}_3(x) & = & x^3-2x \\
         \widetilde{B}_4(x) & = & x^4-3x^2+1 \\
         \widetilde{B}_5(x) & = & x^5 -4x^3 + 3x
       \end{array}
\right.
\end{eqnarray}
with the recurrence (\ref{eq:3.18}).
\begin{eqnarray}\label{eq:3.18}
\widetilde{B}_n(x) = x\widetilde{B}_{n-1}(x)- \widetilde{B}_{n-2}(x),\quad n\geqslant 2.
\end{eqnarray}
Note that this recurrence is the same as that for the Boubaker polynomials but with different initial conditions. In fact the relation between both families of polynomials is given by
\begin{eqnarray}\label{eq:3.19}
T(1+3x^2|1+x^2) = T(1+3x^2|1)T(1|1+x^2).
\end{eqnarray}
Then, finally:
\begin{eqnarray}\label{eq:3.20}
B_n(x) = x\widetilde{B}_{n-1}(x) + 3\widetilde{B}_{n-2}(x),\quad n\geqslant2.
\end{eqnarray}

\section{Fermat-linked expressions}

Using inversion of Riordan matrices we can get $\widetilde{B}_n(x)$ each as combinations of Boubaker polynomials.

\begin{remark}
Comparing the recurrence (\ref{eq:3.20}) with the one of the
Chebyshev polynomials of the second kind, we can obtain an explicit
expression of the new polynomials defined by
(\ref{eq:3.8}-\ref{eq:3.9})
\begin{eqnarray}\label{eq:4.1}
B_n(x) = \frac{\sin((n+1)t)}{\sin t}, \quad x = 2\cos t, \quad n =0,
1,2,\ldots.
\end{eqnarray}
In another word, the new polynomial  is the scaled Chebyshev
polynomial $U_n(x)$ of the second kind, since the relation between
the two polynomials is related as:
\begin{eqnarray}\label{eq:4.2}
B_n(2x) = U_n(x), \quad n=0,1,2,\ldots.
\end{eqnarray}
\end{remark}

\begin{remark}
By using (\ref{eq:4.1}) or (\ref{eq:4.2}), we can obtain some other
relations. In fact Fermat polynomials are obtained by setting
$p(x)=3x$ and $q(x)=-2$ in the Lucas polynomial sequence, defined by
(\ref{eq:4.3}).
\begin{eqnarray}\label{eq:4.3}
F_n(x) = p(x)F_{n-1}(x)+q(x)F_{n-2}(x).
\end{eqnarray}
As A. Luzon and M. A. Moron \cite{Luzon1,Luzon2,Luzon3,Luzon4}
demonstrated, through the associated Riordan matrix:
\begin{eqnarray}\label{eq:4.4}
\left[
  \begin{array}{ccccccccc}
    \frac13 &  &  &  &  &  &  &  &  \\
    0 & 1 &  &  &  &  &  &  &  \\
    0 & 0 & 3 &  &  &  &  &  &  \\
    0 & -2 & 0 & 9 &  &  &  &  &  \\
    0 & 0 & -12 & 0 & 27 &  &  &  &  \\
    0 & 4 & 0 & -54 & 0 & 81 &  &  &  \\
    0 & 0 & 36 & 0 & -216 & 0 & 243 &  &  \\
    0 & -8 & 0 & 216 & 0 & -810 & 0 & 729 &  \\
    \vdots & \vdots & \vdots & \vdots & \vdots & \vdots & \vdots & \vdots & \ddots \\
  \end{array}
\right]
\end{eqnarray}
that
\begin{eqnarray}\label{eq:4.5}
\left\{\begin{array}{rcl}
         F_1(x) & = & 1 \\
         F_2(x) & = & 3x \\
         F_3(x) & = & 9x^2-2 \\
         F_4(x) & = & 27x^3-12x \\
         \cdots & \cdots & \cdots
       \end{array}
 \right.
\end{eqnarray}
and
\begin{eqnarray}\label{eq:4.6}
F_x(x) = \Pa{\sqrt{2}}^n U_n\Pa{\frac{3x}{2\sqrt{2}}}
\end{eqnarray}
\end{remark}

\begin{thrm}
Let $(R,+,o)$ be a commutative ring, $(D,+,o)$ be an integral domain
such that $D$ is a subring of $R$ whose zero is $0_D$ and whose
unity is $1_D$, $X\in R$ be transcendental over $D$, $D[X]$ be the
ring of polynomials forms in $X$ over $D$, and finally denote
Boubaker polynomials and Fermat polynomials as $B_n(x)$ and $F_n(x)$
,respectively, as polynomials contained in $D[X]$, then:
\begin{eqnarray}\label{eq:4.7}
B_n(x) =\frac1{(\sqrt{2})^n}F_n\Pa{\frac{2\sqrt{2}x}3} +
\frac1{(\sqrt{2})^{n-2}}F_{n-2}\Pa{\frac{2\sqrt{2}x}3};\quad n =
0,1,2,\ldots
\end{eqnarray}
\end{thrm}

\begin{proof}
Riordan matrices for Boubaker polynomials and Fermat polynomials
(see \cite{Luzon1,Luzon2,Luzon3,Luzon4}) are respectively:
\begin{eqnarray}\label{eq:4.8}
\sum^{+\infty}_{n=0}B_n(x)t^n =
(1+3x^2|1+x^2)\Pa{\frac1{1-xt}},\quad \sum^{+\infty}_{n=0}F_n(x)t^n
=\Pa{\frac13|\frac{1+x^2}3}.
\end{eqnarray}
Let's expand the inverse Riordan arrays:
\begin{eqnarray}\label{eq:4.9}
T(1+3x^2|1+x^2) = T(1+3x^2|1)T\Pa{\frac12|\frac{1+x^2}2}T(2|2),
\end{eqnarray}
which gives
\begin{eqnarray}\label{eq:4.10}
T(1+3x^2|1+x^2) =
T(1+3x^2|1)T(1|\sqrt{2})T\Pa{\frac13|\frac{1+x^2}3}T(3|\frac3{\sqrt{2}}).
\end{eqnarray}
By identifying Riordan matrix for Fermat polynomials in the right
term of Eq. (\ref{eq:4.10}), the desired equality holds.
\end{proof}

Expressions (\ref{eq:4.2}) and (\ref{eq:4.7}) are very useful for
developing the already proposed Boubaker polynomials Expansion
Scheme (BPES).

\section{Conclusion}

The Boubaker polynomials have been  investigated. Using y Riordan
matrices analysis, a sequence of relations outlining the relations
with Chebyshev and Fermat polynomials have been obtained as guides
to further studies. The obtained expression are a meaningful supply
to recent applied physics studies
\cite{Mahmoud1,Lazzez,Fridjine1,Khelia,Mahmoud2,Dada,Rahmanov1,Rahmanov2,Tabatabaei,Fridjine2,Belhadj1,Belhadj2,Zhang}
using the Boubaker polynomials Expansion Scheme (BPES).


\end{document}